# Linear Noise Approximation Assisted Bayesian Inference on Mechanistic Model of Partially Observed Stochastic Reaction Network

### v1

@@ 4 BAYESIAN ANALYSIS AND Algorithm Development
$p\left(\boldsymbol{\theta},\boldsymbol{\sigma}|\mathcal{D}_{M}\right)$ $\propto p\left(\boldsymbol{\theta}\right)p\left(\boldsymbol{\sigma}\right)\exp\left\{-\frac{1}{2}\left[M|\boldsymbol{J}_{0}|\log(2\pi)+\log\left|\boldsymbol{G}_{0}\boldsymbol{\Psi}(t_{0})\boldsymbol{G}_{0}^{\top}+\boldsymbol{\Sigma}_{0}\right|+\right.\right.$
$\quad\left.\left(\boldsymbol{y}_{0}(t_{0})-\boldsymbol{G}_{0}\left\{\bar{\boldsymbol{s}}(t_{0})+\boldsymbol{\varphi}(t_{0})\right\}\right)^{\top}\left(\boldsymbol{G}_{0}\boldsymbol{\Psi}(t_{0})\boldsymbol{G}_{0}^{\top}+\boldsymbol{\Sigma}_{0}\right)^{-1}\left(\boldsymbol{y}_{0}(t_{0})-\boldsymbol{G}_{0}\left\{\bar{\boldsymbol{s}}(t_{0})+\boldsymbol{\varphi}(t_{0})\right\}\right)\right]-$
$\quad\frac{1}{2}\sum_{h=0}^{H-1}\left[M|\boldsymbol{J}_{h+1}|\log(2\pi)+\log\left|\boldsymbol{G}_{h+1}\boldsymbol{\Psi}(t_{h+1})\boldsymbol{G}_{h+1}^{\top}+\boldsymbol{\Sigma}_{h+1}\right|+\left(\boldsymbol{y}_{h+1}(t_{h+1})-\boldsymbol{G}_{h+1}\bar{\boldsymbol{s}}(t_{h+1})\right)^{\top}\right.$
$\quad\left.\left.\left(\boldsymbol{G}_{h+1}\boldsymbol{\Psi}(t_{h+1})\boldsymbol{G}_{h+1}^{\top}+\boldsymbol{\Sigma}_{h+1}\right)^{-1}\left(\boldsymbol{y}_{h+1}(t_{h+1})-\boldsymbol{G}_{h+1}\bar{\boldsymbol{s}}(t_{h+1})\right)\right]\right\},$ (23)
@@
To numerically solve Equation (24) and generate posterior samples from $p(\boldsymbol{\eta}|\mathcal{D}_{M})$, the Euler-Maruyama approximation is used to obtain the discretized Langevin diffusion with a step size $\Delta\tau>0$,
@@
To correct the bias in the stationary distribution induced by the discretization used in the update rule (25) for the posterior search, a Metropolis-Hastings step is incorporated for simulating the Langevin diffusion (24). Specifically, we consider the update rule (25) and define a proposal distribution to generate a new MCMC posterior sample $\tilde{\boldsymbol{\eta}}(\tau+1)$,
@@ 5 EMPIRICAL STUDY
We assess the performance of the proposed MALA joint posterior sampler under model uncertainty induced with the different data sizes, i.e., $H=4,8,16$ ($\Delta t=20,10,5$ seconds correspondingly). Its performance is compared with classic Metropolis-Hastings algorithm (M-H) in terms of: (1) convergence speed, and (2) convergence results on model parameters. Notice that both two algorithms are based on the likelihood approximated by the Bayesian updating LNA metamodel. And since the support of the parameters is the positive space, we need to use the logarithmic function to transform it to the real space at first. That is, we set $\log(\boldsymbol{\eta})=\left(\log(\theta_{1}),\log(\theta_{2}),\log(\theta_{3}),\log(\sigma)\right)^{\top}$ and apply both algorithms to $\log(\boldsymbol{\eta})$. For Bayesian updating LNA metamodel, we set $\bar{\boldsymbol{s}}(t_{0})+\boldsymbol{\varphi}(t_{0})=\left(50,40,60,10\right)^{\top}$, $\boldsymbol{\Psi}(t_{0})$ to a 4-by-4 identity matrix, and $I_{h}=2000,1000,500$ for $\Delta t=20,10,5$ respectively to make a small $\Delta z_{h}=0.01$ for any $h=0,1,\ldots,H$. The priors of the parameters are set as $\theta_{k}\sim U(0,1)$ for $k=1,2,3$, and $\sigma\sim U(0,25)$, to consider a difficult situation without strong prior information. The results are estimated based on 10 macro-replications.
@@
First, we compare the convergence speed of the MALA and M-H approaches. For MALA, we set the step size $\Delta\tau=0.001$. Correspondingly, to show that MALA improves the mixing of MCMC. For M-H, we set its proposal distribution to be Gaussian distributed with the current sample as mean and diag$\{2\Delta\tau\}=$ diag$\{0.002\}$ as covariance. Figure 2 shows the mean convergence trends of the two algorithms for the three log-kinetic rate parameters (with 95% confidence intervals across 10 macro-replications) under the data size $H=16\ (\Delta t=5)$. The black line represents the true log-parameters. By comparing the widths of the confidence intervals as iterations progress, we observe that MALA shows a significant improvement in the convergence speed of the log-kinetic rate parameters inference over M-H, demonstrating that by sufficiently leveraging on the likelihood information provided by a moderate size of observations, MALA posterior sampling converges quickly towards the true log-parameters.
@@
Figure 2: The convergence trends of MALA and Metropolis-Hastings algorithms (with 95% confidence intervals) when the data size $H=16\ (\Delta t=5)$.
@@
Then, we study the root mean square error (RMSE) of model parameter estimation to assess the convergence results of the MALA and M-H posterior sampling algorithms under three different data sizes. Basically, the RMSE measures the differences between the true and the estimated log-parameters based on $B$ posterior samples, i.e., RMSE $=\sqrt{\frac{1}{B}\sum_{b=1}^{B}|\log(\eta_{l}^{\rm true})-\{\log(\eta_{l})\}^{(b)}|^{2}}$ for $l=1,2,3,4$. We set the initial warm-up length $T_{0}=10000$ to reduce the initial bias, the posterior sample size $B=100$, and an appropriate integer $\delta=10$ to reduce sample correlation for both algorithms. We summarize the 95% confidence intervals obtained by using 10 macro-replications of the RMSEs for the four log-parameters inferred by the two algorithms in Table 1. As it shows, for the four log-parameters and the two algorithms except $\log(\theta_{3})$ inferred by MALA under the data size $H=8$ and $H=16$, the RMSEs decrease significantly as the data size increases, demonstrating the major benefit induced by the Bayesian updating LNA metamodel. That means it refines the approximation of the likelihood as more observations available to update the LNA model. The exception of $\log(\theta_{3})$ is due to that, $\theta_{3}$ is the kinetic rate parameter associated with the Reaction 3, whose reactant (i.e., Complex) is the only observable component of this system, the data size of $H=8$ is thus sufficient to provide relatively accurate likelihood information. With the help of MALA, $\log(\theta_{3})$ converges close to the true value when $H=8$. Additionally, for $\log(\theta_{1})$ and $\log(\theta_{2})$ under all data sizes, MALA performs better than M-H in terms of both RMSEs and their confidence half-widths, meaning that MALA converges faster than M-H; while for $\log(\theta_{3})$ and $\log(\sigma)$, the performance of the two algorithms is similar. This is because both two algorithms have converged before $T_{0}=10000$.
Table 1: The RMSEs between the estimated and the true log-parameters (with 95% confidence intervals).
Algorithm MALA Metropolis-Hastings (M-H)
Data size $H=4$ $H=8$ $H=16$ $H=4$ $H=8$ $H=16$
$(\Delta t=20)$ $(\Delta t=10)$ $(\Delta t=5)$ $(\Delta t=20)$ $(\Delta t=10)$ $(\Delta t=5)$
$\log(\theta_{1})$ $1.79\pm 0.73$ $1.27\pm 0.92$ $0.48\pm 0.08$ $2.56\pm 1.09$ $1.80\pm 1.07$ $1.48\pm 1.37$
$\log(\theta_{2})$ $2.76\pm 1.37$ $2.12\pm 1.31$ $1.32\pm 0.58$ $3.29\pm 1.82$ $2.87\pm 1.49$ $2.50\pm 1.61$
$\log(\theta_{3})$ $1.73\pm 2.02$ $0.25\pm 0.03$ $0.28\pm 0.03$ $1.66\pm 2.12$ $1.15\pm 2.02$ $0.24\pm 0.03$
$\log(\sigma)$ $1.81\pm 0.93$ $1.02\pm 0.56$ $0.92\pm 0.59$ $1.59\pm 0.70$ $0.98\pm 0.43$ $0.80\pm 0.29$
@@ 6 Conclusion
Bayesian inference on partially observed SRN plays a critical role for multi-scale bioprocess mechanistic modeling. To tackle the critical challenges of biomanufacturing processes, including high complexity, high inherent stochasticity, and very limited observations on partially observed state with measurement errors, we propose an interpretable Bayesian updating LNA metamodel to approximate the likelihood of heterogeneous online and offline measures, accounting for the structure information of the enzymatic SRN mechanistic model. Then, we develop a MALA sampling algorithm that utilizes the information from the derived likelihood and more efficiently generates posterior samples. The empirical study shows that our proposed LNA assisted Bayesian inference approach has a promising performance, demonstrating its potential to benefit bioprocess mechanisms online learning and digital twin development.

### v2

@@ 4 BAYESIAN ANALYSIS AND Algorithm Development
$p\left(\boldsymbol{\theta},\boldsymbol{\sigma}|\mathcal{D}_{M}\right)$ $\propto p\left(\boldsymbol{\theta}\right)p\left(\boldsymbol{\sigma}\right)\exp\left\{-\frac{1}{2}\left[M|\boldsymbol{J}_{0}|\log(2\pi)+\log\left|\boldsymbol{G}_{0}\boldsymbol{\Psi}(t_{0})\boldsymbol{G}_{0}^{\top}+\boldsymbol{\Sigma}_{0}\right|\right.\right.$
$\quad\left.+\left(\boldsymbol{y}_{0}(t_{0})-\boldsymbol{G}_{0}\left\{\bar{\boldsymbol{s}}(t_{0})+\boldsymbol{\varphi}(t_{0})\right\}\right)^{\top}\left(\boldsymbol{G}_{0}\boldsymbol{\Psi}(t_{0})\boldsymbol{G}_{0}^{\top}+\boldsymbol{\Sigma}_{0}\right)^{-1}\left(\boldsymbol{y}_{0}(t_{0})-\boldsymbol{G}_{0}\left\{\bar{\boldsymbol{s}}(t_{0})+\boldsymbol{\varphi}(t_{0})\right\}\right)\right]$
$\quad-\frac{1}{2}\sum_{h=0}^{H-1}\left[M|\boldsymbol{J}_{h+1}|\log(2\pi)+\log\left|\boldsymbol{G}_{h+1}\boldsymbol{\Psi}(t_{h+1})\boldsymbol{G}_{h+1}^{\top}+\boldsymbol{\Sigma}_{h+1}\right|\right.$
$\quad\left.\left.+\left(\boldsymbol{y}_{h+1}(t_{h+1})-\boldsymbol{G}_{h+1}\bar{\boldsymbol{s}}(t_{h+1})\right)^{\top}\left(\boldsymbol{G}_{h+1}\boldsymbol{\Psi}(t_{h+1})\boldsymbol{G}_{h+1}^{\top}+\boldsymbol{\Sigma}_{h+1}\right)^{-1}\left(\boldsymbol{y}_{h+1}(t_{h+1})-\boldsymbol{G}_{h+1}\bar{\boldsymbol{s}}(t_{h+1})\right)\right]\right\},$ (23)
@@
To numerically solve Equation (24) and generate posterior samples from $p(\boldsymbol{\eta}|\mathcal{D}_{M})$, the Euler-Maruyama approximation [13] is used to obtain the discretized Langevin diffusion with a step size $\Delta\tau>0$,
@@
To correct the bias in the stationary distribution induced by the discretization used in the update rule (25), a Metropolis-Hastings step is incorporated for simulating the Langevin diffusion (24). Specifically, we consider the update rule (25) and define a proposal distribution to generate a new MCMC posterior sample $\tilde{\boldsymbol{\eta}}(\tau+1)$,
@@ 5 EMPIRICAL STUDY
We assess the performance of the proposed joint posterior sampler under model uncertainty induced with the different data sizes, i.e., $H=4,8,16$ ($\Delta t=20,10,5$ seconds correspondingly). To study the effect of additional gradient information and Bayesian updating step, the MALA with Bayesian updating LNA is compared to other candidate approaches, including (1) classic Metropolis-Hastings algorithm (M-H) with Bayesian updating LNA, and (2) MALA with original LNA (without Bayesian update), in terms of convergence behavior of posterior sampling. Since the support of the parameters is the positive space, we first need to use the logarithmic function to transform it to the real space. That is, we set $\log(\boldsymbol{\eta})=\left(\log(\theta_{1}),\log(\theta_{2}),\log(\theta_{3}),\log(\sigma)\right)^{\top}$ and apply three algorithms to $\log(\boldsymbol{\eta})$. For both LNA metamodels, we set $\bar{\boldsymbol{s}}(t_{0})+\boldsymbol{\varphi}(t_{0})=\left(50,40,60,10\right)^{\top}$, $\boldsymbol{\Psi}(t_{0})$ as a 4-by-4 identity matrix, and $I_{h}=2000,1000,500$ for $\Delta t=20,10,5$ respectively to make a small $\Delta z_{h}=0.01$ for any $h=0,1,\ldots,H$. The priors of the parameters are set as $\theta_{k}\sim U(0,1)$ for $k=1,2,3$, and $\sigma\sim U(0,25)$, to consider a difficult situation without strong prior information. The results are estimated based on 10 macro-replications.
@@
First, we compare the convergence speed of three algorithms. For MALA with Bayesian updating LNA and with original LNA (without Bayesian update), we set the step size $\Delta\tau=0.001$. Correspondingly, to show that MALA improves the mixing of MCMC, for M-H with Bayesian updating LNA, we set its proposal distribution to be Gaussian distributed with the current sample as mean and diag$\{2\Delta\tau\}=$ diag$\{0.002\}$ as covariance. Figure 2 shows the mean convergence trends of the three algorithms for the three log-kinetic rate parameters (with 95% confidence intervals (CIs) across 10 macro-replications) under the data size $H=16\ (\Delta t=5)$. The black line represents the true log-parameters. By comparing the widths of the CIs as iterations progress, we observe that MALA shows a significant improvement in the convergence speed of the log-kinetic rate parameters inference over M-H, while the Bayesian updating step reduces the approximation error accumulation over time of original LNA, providing a more accurate likelihood approximation. It demonstrates that by sufficiently leveraging on the likelihood and its gradient information provided by a moderate size of observations, MALA posterior sampling based on the likelihood approximated by the Bayesian updating LNA metamodel converges quickly towards the true log-parameters.
@@
Figure 2: The convergence trends of (1) MALA with original LNA, (2) MALA with Bayesian updating LNA, and (3) M-H with Bayesian updating LNA (with 95% CIs) when the data size $H=16\ (\Delta t=5)$.
@@
Then, we study the root mean square error (RMSE) of model parameter estimation to assess the convergence results of three algorithms under three different data sizes. Basically, the RMSE measures the differences between the true and the estimated log-parameters based on $B$ posterior samples, i.e., RMSE $=\sqrt{\frac{1}{B}\sum_{b=1}^{B}|\log(\eta_{l}^{\rm true})-\{\log(\eta_{l})\}^{(b)}|^{2}}$ for $l=1,2,3,4$. We set the initial warm-up length $T_{0}=10000$ to reduce the initial bias, the posterior sample size $B=100$, and an appropriate integer $\delta=10$ to reduce sample correlation for three algorithms. We summarize the 95% CIs obtained by using 10 macro-replications of the RMSEs for the four log-parameters inferred by the three algorithms in Table 1. As it shows, for the four log-parameters except $\log(\theta_{3})$ inferred under the data size $H=8,16$, the RMSEs of MALA with Bayesian updating LNA decrease more significantly than those of MALA with original LNA as the data size increases, demonstrating the major benefit induced by the Bayesian updating LNA metamodel. That means it refines the approximation of the likelihood as more observations available to update the original LNA model. The exception of $\log(\theta_{3})$ is due to that, $\theta_{3}$ is the kinetic rate parameter associated with the Reaction 3, whose reactant (i.e., Complex) is the only observable component of this system, the data size of $H=16$ is thus sufficient to provide relatively accurate likelihood information based on the original LNA metamodel, while the Bayesian updating step improves accuracy of likelihood approximation when the data size $H=8$. With the help of MALA, $\log(\theta_{3})$ converges close to the true value when $H=8,16$. Additionally, for $\log(\theta_{1})$ and $\log(\theta_{2})$ under all data sizes, based on the likelihood approximated by the Bayesian updating LNA metamodel, MALA performs better than M-H in terms of both RMSEs and their confidence half-widths, meaning that MALA converges faster than M-H; while for $\log(\theta_{3})$ and $\log(\sigma)$, the performance of the two algorithms is similar. This is because both two algorithms have converged before $T_{0}=10000$.
Table 1: The RMSEs between the estimated and the true log-parameters (with 95% CIs).
MALA with Bayesian updating LNA M-H with Bayesian updating LNA MALA with original LNA
Data $H=4$ $H=8$ $H=16$ $H=4$ $H=8$ $H=16$ $H=4$ $H=8$ $H=16$
size $(\Delta t=20)$ $(\Delta t=10)$ $(\Delta t=5)$ $(\Delta t=20)$ $(\Delta t=10)$ $(\Delta t=5)$ $(\Delta t=20)$ $(\Delta t=10)$ $(\Delta t=5)$
$\log(\theta_{1})$ $1.79\pm 0.73$ $1.27\pm 0.92$ $0.48\pm 0.08$ $2.56\pm 1.09$ $1.80\pm 1.07$ $1.48\pm 1.37$ $1.82\pm 0.74$ $1.72\pm 0.81$ $1.41\pm 0.94$
$\log(\theta_{2})$ $2.76\pm 1.37$ $2.12\pm 1.31$ $1.32\pm 0.58$ $3.29\pm 1.82$ $2.87\pm 1.49$ $2.50\pm 1.61$ $2.80\pm 1.38$ $2.73\pm 1.36$ $2.63\pm 1.19$
$\log(\theta_{3})$ $1.73\pm 2.02$ $0.25\pm 0.03$ $0.28\pm 0.03$ $1.66\pm 2.12$ $1.15\pm 2.02$ $0.24\pm 0.03$ $1.71\pm 2.02$ $0.69\pm 0.83$ $0.17\pm 0.04$
$\log(\sigma)$ $1.81\pm 0.93$ $1.02\pm 0.56$ $0.92\pm 0.59$ $1.59\pm 0.70$ $0.98\pm 0.43$ $0.80\pm 0.29$ $1.76\pm 0.93$ $1.30\pm 0.62$ $1.05\pm 0.73$
@@ 6 Conclusion
Bayesian inference on partially observed SRN plays a critical role for multi-scale bioprocess mechanism learning. To tackle the critical challenges of biomanufacturing processes, including high complexity, high inherent stochasticity, and very limited and sparse observations on partially observed state with measurement errors, we propose an interpretable Bayesian updating LNA metamodel to approximate the likelihood of heterogeneous online and offline measures, accounting for the structure information of the enzymatic SRN mechanistic model. Then, we develop a MALA sampling algorithm that utilizes the information from the derived likelihood and more efficiently generates posterior samples. The empirical study shows that our proposed LNA assisted Bayesian inference approach has a promising performance, demonstrating its potential to benefit bioprocess mechanisms online learning and digital twin development.
Acknowledgements
The authors thank Prof. Cheng Li from the Department of Statistics and Data Science, National University of Singapore for providing the discussion of Metropolis-adjusted Langevin algorithm (MALA).